OpenStudy (anonymous):

What is the area of a circle with a radius of 4 meters? Use 22/7 for pie .

OpenStudy (anonymous):

Area of a circle is given by \(\pi r^2 \), where \( r\) is the radius. So, \(\frac {22}7 \times 4^2 = \frac{352}{7} \) sq mt

OpenStudy (anonymous):

\( = 50.2857\) sq mt

OpenStudy (anonymous):

@FoolForMath, do you guys use the abbreviation \(\text{mt}\) rather than \(\text{m}?\) In the states, it's always \(\text{m}.\) Just curious.

OpenStudy (anonymous):

No we use 'm' typically but I have seen the usage of 'mt' :)

OpenStudy (anonymous):

Interesting. You learn something new everyday. :P